### v1

@@ Abstract
The matrix code equivalence problem consists, given two matrix spaces $\mathcal{C},\mathcal{D}\subset\mathbb{F}_{q}^{m\times n}$ of dimension $k$, in finding invertible matrices $P\in\operatorname{GL}_{m}(\mathbb{F}_{q})$ and $Q\in\operatorname{GL}_{n}(\mathbb{F}_{q})$ such that $\mathcal{D}=P\mathcal{C}Q^{-1}$. Recent signature schemes such as MEDS and ALTEQ relate their security to the hardness of this problem. Naranayan et. al. recently published an algorithm solving this problem in the case $k=n=m$ in $\widetilde{\mathcal{O}}(q^{\frac{k}{2}})$ operations. We present a different algorithm which solves the problem in the general case. Our approach consists in reducing the problem to the matrix code conjugacy problem, i.e. the case $P=Q$. For the latter problem, similarly to the permutation code equivalence problem in Hamming metric, a natural invariant based on the Hull of the code can be used. Next, the equivalence of codes can be deduced using a usual list collision argument. For $k=m=n$, our algorithm achieves the same complexity as in the aforementioned reference. However, it extends to a much broader range of parameters.
@@ Introduction
In the last decades so-called equivalence problems have frequently been used for cryptographic applications. The first examples probably come from multivariate cryptography with Matsumoto Imai [19] or HFE [22] schemes whose security relies on the hardness of the polynomial isomorphism problem: deciding whether two spaces of polynomials are equivalent with respect to a linear or affine change of variables.
In the recent years, we observed an intensification of this trend but also a diversification of the equivalence problems used for cryptography. In particular, in NIST’s recent on-ramp call for signature [^1] https://csrc.nist.gov/Projects/pqc-dig-sig/round-1-additional-signatures, many signature schemes involve equivalence problems which are not the polynomial isomorphism one. For instance, Hawk’s [5] security rests among others on the hardness of the Lattice Isomorphism Problem (LIP), LESS [1] rests on the monomial equivalence of Hamming metric codes, MEDS [9] on the matrix code equivalence and ALTEQ [4] on the equivalence of alternate trilinear forms. The latter problem (equivalence of alternate trilinear forms) is actually a sub-case of the former: the matrix code equivalence problem, which is the purpose of the present article. The other way around the two problems are proved to be polynomially equivalent [14, Prop. 8.3].
Given two matrix spaces $\mathcal{C},\mathcal{D}\subset\mathbb{F}_{q}^{m\times n}$, the matrix code equivalence problem consists in deciding whether there exists $P\in\operatorname{GL}_{m}(\mathbb{F}_{q})$ and $Q\in\operatorname{GL}_{n}(\mathbb{F}_{q})$ such that $\mathcal{C}=P\mathcal{D}Q^{-1}$. The search version of the problem asks to return, if exists, a pair $P,Q$ providing the equivalence. If its use for cryptography is rather new, this problem is known for a long time in algebraic complexity theory where it is usually formulated in an equivalent way as the $3$–tensor isomorphism problem. This problem is assumed to be hard and is in particular known to be at least as hard as the monomial code equivalence problem (see [14] or [11]).
@@
In this article, we propose a new algorithm for solving the matrix code equivalence problem, or equivalently the $3$–tensor isomorphism problem. Given equivalent $k$–dimensional $m\times n$ matrix spaces with entries in $\mathbb{F}_{q}$, we are able to solve the search equivalence problem in a complexity $\widetilde{\mathcal{O}}(q^{s-\frac{m-1}{2}})$ where $s=\min\{k,mn-k\}$. Note in particular that in the specific case $k=m=n$ which is the one that is used in the parameters of MEDS and ALTEQ, we achieve the complexity $\widetilde{\mathcal{O}}(q^{\frac{k}{2}})$ which is the one achieved by Narayanan, Qiao and Tang [20]. However,
@@
MEDS and ALTEQ schemes [9, 4] were both submitted to NIST’s on-ramp call for digital signatures. Before, ALTEQ’s and MEDS’ specifications were respectively presented in the articles [29] and [10]. In [3], Beullens proposes a new algorithm solving the trilinear form equivalence problem, harming the proposed parameters for ALTEQ. More recently, Narayanan, Qiao and Tang [20] proposed an algorithm solving the same problem but also the matrix code equivalence problem in the case of $k$–dimensional spaces of $k\times k$ matrices. Their approach combines a collision list argument with a nice algebraic invariant and achieves a complexity in $\widetilde{\mathcal{O}}(q^{\frac{k}{2}})$.
Outline of the article
In Section 1, we introduce the matrix code equivalence problem as well as some related problems and algorithms solving them. In Section 2, we first state some key observations at the heart of our algorithm, then give a brief presentation of the algorithm as well as its complexity. The algorithm itself can be divided into three steps. The crucial one is the reduction to a specific instance of the matrix code conjugacy problem, where conjugate one-dimensional subspaces inside the two codes are given. This reduction is presented in Section 3. We then solve the conjugacy problem in question using an algorithm presented in Section 4. Finally, we explain in Section 5 how to deduce a solution to the initial matrix code equivalence problem from the solution to the conjugacy problem found in the previous step.
Acknowledgements
The second author is funded by Inria and the French Cybersecurity Agency (ANSSI). This work benefited from the financial support of the French government and the Agence Nationale de la Recherche (ANR) through the Plan France 2030 via the project ANR-22-PETQ-0008. The authors are part of a collaborative research project Barracuda with reference ANR-21-CE39-0009-BARRACUDA. The first author is partially funded by Horizon-Europe MSCA-DN project Encode.
@@ 1 The matrix code equivalence problem
so any algorithm solving the case $m\leqslant n$ can also be used, after transposing the whole problem, to solve the case $n\leqslant m$. In the remainder of this article, we will always suppose that $m\leqslant n$.
@@ 1.1.1 The trilinear forms equivalence problem
$(Px,Qy,Rz)=g(x,y,z).$
@@
$f$ $\colon(x,y,z)\mapsto\sum_{i,j,k}c_{ijk}x_{i}y_{j}z_{k}$
$g$ $\colon(x,y,z)\mapsto\sum_{i,j,k}d_{ijk}x_{i}y_{j}z_{k}.$
@@ 1.1.2 Alternate trilinear form equivalence problem
The alternate trilinear forms equivalence problem $\mathsf{TFE}_{m,n,k}$ is the following. Given two alternate trilinear forms $f,g\colon\mathbb{F}_{q}^{m}\times\mathbb{F}_{q}^{m}\times\mathbb{F}_{q}^{k}\to\mathbb{F}_{q}$, find a matrix $P\in\operatorname{GL}_{n}(\mathbb{F}_{q})$ such that for any $x,y,z\in\mathbb{F}_{q}^{m}\times\mathbb{F}_{q}^{n}\times\mathbb{F}_{q}^{k}$,
@@
Rewriting an instance of $\mathsf{MCE}$ as an instance of $\mathsf{TFE}$ shows that the problem is symmetric in the three parameters $m,n,k$. In particular, we may choose to permute $m,n,k$ as we like in our algorithm in order to minimize its complexity. Moreover, as we will explain in Lemma 2, we may also switch $k$ for $mn-k$.
@@ 1.2 Related works on attacks on $\mathsf{MCE}$ and $\mathsf{ATFE}$
For the design of MEDS [10], the authors consider a graph-search based approach inspired from the works of Bouillaguet, Fouque and Véber [6] on the polynomial isomorphism problem for spaces of quadratic forms. This approach leads to an attack of complexity $\widetilde{\mathcal{O}}(q^{\frac{2}{3}(m+n)})$. Also, they consider the possibility of algebraic modelling which turns out to be harder than for $\mathsf{ATFE}$ since the unknown correspond to a pair of matrices $(P,Q)$ instead of a single one. They also study a “Leon–like” approach, a reference to Leon’s algorithm [17] for determining code equivalence that consists in harvesting minimum weight codewords to determine the code equivalence. When transposed to the matrix code setting, the Hamming weight is replaced by the rank and such approach is nothing but the aforementioned MinRank based technique, which, following a recent result from Beullens [2] on Hamming metric code equivalence, can be combined with a collision search technique. MEDS’ parameter selection rests on the complexity of both algebraic attacks and Leon–like ones.
@@
First, Beullens proposed a graph–search–based technique to solve $\mathsf{ATFE}$. His attack turns out to be particularly efficient for low values of $n$. For instance, for $n$ odd, he could achieve a complexity in $\mathcal{O}(q^{{(n-5)}/{2}}n^{11}+q^{n-7}n^{6})$. This permitted to identify weak keys in [29].
@@
In [20], the authors introduced a new algorithm solving both $\mathsf{ATFE}$ and $\mathsf{MCE}$ in the specific case $k=m=n$. We conclude this section by sketching the principle of this algorithm in order to point out the need for being in the cubic case $k=m=n$.
@@
The idea of the algorithm consists first in guessing a pair $(x_{1},x_{1}^{\prime})\in{(\mathbb{F}_{q}^{n})}^{2}$ such that $x_{1}^{\prime}=Px_{1}$ and $f(x_{1},\cdot,\cdot)$ is a bilinear form of rank $n-1$. Next, due to the rank constraint, from $x_{1}$ can be deduced a unique $y_{1}$ (up to scalar multiplication) such that $f(x_{1},y_{1},\cdot)\equiv 0$. Similarly, they deduce a $z_{1}$ such that $f(\cdot,y_{1},z_{1})\equiv 0$, and an $x_{2}$ such that $f(x_{2},\cdot,z_{1})\equiv 0$ and so on. By this manner, they construct 3 sequences $x_{1},\dots,x_{n}$, $y_{1},\dots,y_{n}$ and $z_{1},\dots,z_{n}$ for $f$ and similarly construct $x_{1}^{\prime},\cdots,x_{n}^{\prime}$, $y_{1}^{\prime},\dots,y_{n}^{\prime}$ and $z_{1}^{\prime},\dots,z_{n}^{\prime}$ for $f^{\prime}$. Stacking these vectors as columns of $n\times n$ matrices, we get 6 matrices $X,Y,Z,X^{\prime},Y^{\prime}$ and $Z^{\prime}$ that will be invertible with a high probability.
@@
In view of this observation, the algorithm solving $\mathsf{MCE}$ consists in a collision search between two dictionaries. The first one collects pairs $(f_{x_{1}},x_{1})$, the left–hand term being the search key and the right–hand one being the corresponding value, and the second one collects pairs $(g_{x_{1}^{\prime}},x_{1}^{\prime})$. Once such a collision is found, which requires $O(q^{n/2})$ trials, determining the equivalence becomes easy (see [20] for further details).
@@
It should be emphasized that the crux of their algorithm rests on the unique possibility (up to scalar multiplication) of passing from $x_{i}$ to $y_{i}$, from $y_{i}$ to $z_{i}$ and from $z_{i}$ to $x_{i+1}$. Such a technique is possible only because at each step, the corresponding bilinear form is represented by a rank $n-1$ matrix of size $n\times n$. Hence, their approach strongly rests on the fact that all the involved matrices have the same sizes, i.e. that they lie in the cubic case $k=m=n$.

### v2

@@ Abstract
The matrix code equivalence problem consists, given two matrix spaces $\mathcal{C},\mathcal{D}\subset\mathbb{F}_{q}^{m\times n}$ of dimension $k$, in finding invertible matrices $P\in\mathrm{GL}_{m}(\mathbb{F}_{q})$ and $Q\in\mathrm{GL}_{n}(\mathbb{F}_{q})$ such that $\mathcal{D}=P\mathcal{C}Q^{-1}$. Recent signature schemes such as MEDS and ALTEQ relate their security to the hardness of this problem. Recent works by Narayanan, Qiao and Tang on the one hand and by Ran and Samardjiska on the other hand tackle this problem. The former is restricted to the “cubic” case $k=m=n$ and succeeds in $\widetilde{\mathcal{O}}(q^{\frac{k}{2}})$ operations. The latter is an algebraic attack on the general problem whose complexity is not fully understood and which succeeds only on $\mathcal{O}(1/q)$ instances. We present a novel algorithm which solves the problem in the general case. Our approach consists in reducing the problem to the matrix code conjugacy problem, i.e. the case $P=Q$. For the latter problem, similarly to the permutation code equivalence problem in Hamming metric, a natural invariant based on the Hull of the code can be used. Next, the equivalence of codes can be deduced using a usual list collision argument. For $k=m=n$, our algorithm achieves the same time complexity as Narayanan et al. but with a lower space complexity. Moreover, ours extends to a much broader range of parameters.
@@ Introduction
In the last decades, so-called equivalence problems have frequently been used for cryptographic applications. The first examples probably come from multivariate cryptography with the Matsumoto-Imai [19] or HFE [22] schemes, whose security relies on the hardness of the polynomial isomorphism problem: deciding whether two spaces of polynomials are equivalent with respect to a linear or affine change of variables.
In recent years, we observed an intensification of this trend but also a diversification of the equivalence problems used for cryptography. In particular, in NIST’s recent on-ramp call for signature [^1] https://csrc.nist.gov/Projects/pqc-dig-sig/round-1-additional-signatures, many signature schemes involve equivalence problems which are not the polynomial isomorphism one. For instance, Hawk’s [5] security rests among others on the hardness of the Lattice Isomorphism Problem (LIP), LESS [1] rests on the monomial equivalence of Hamming metric codes, MEDS [9] on the matrix code equivalence and ALTEQ [4] on the equivalence of alternate trilinear forms. The latter problem (equivalence of alternate trilinear forms) is in fact a sub-case of the former: the matrix code equivalence problem, which is the purpose of the present article. The two problems are actually proven to be polynomially equivalent [14, Prop. 8.3].
Given two matrix spaces $\mathcal{C},\mathcal{D}\subset\mathbb{F}_{q}^{m\times n}$, the matrix code equivalence problem consists in deciding whether there exists $P\in\operatorname{GL}_{m}(\mathbb{F}_{q})$ and $Q\in\operatorname{GL}_{n}(\mathbb{F}_{q})$ such that $\mathcal{D}=P\mathcal{C}Q^{-1}$. The search version of the problem asks to return, if exists, a pair $(P,Q)$ providing the equivalence. Even though its use for cryptography is rather new, this problem has been known for a long time in algebraic complexity theory where it is usually formulated in an equivalent way as the $3$–tensor isomorphism problem. This problem is assumed to be hard and is in particular known to be at least as hard as the monomial code equivalence problem (see [14] or [11]).
@@
In this article, we present a new algorithm for solving the matrix code equivalence problem, or equivalently the $3$–tensor isomorphism problem. Given equivalent $k$–dimensional $m\times n$ matrix spaces with entries in $\mathbb{F}_{q}$, we are able to solve the search equivalence problem in
$\widetilde{\mathcal{O}}(q^{\max(\frac{k}{2},k-m+2)})$
operations in $\mathbb{F}_{q}$. Note in particular that in the specific case $k=m=n$ which is the one that is used in the parameters of MEDS and ALTEQ, we achieve the time complexity $\widetilde{\mathcal{O}}(q^{\frac{k}{2}})$ which is the one achieved by Narayanan, Qiao and Tang [20]. However,
@@
    in the specific case $k=m=n$, the space complexity of our algorithm is $\mathcal{O}(nq^{\frac{n}{2}-1})$, which is is smaller than that of [20] by a factor $n^{2}$;
  3. 3.
@@
The schemes MEDS and ALTEQ [9, 4] were both submitted to NIST’s on-ramp call for digital signatures. Before, ALTEQ’s and MEDS’ specifications were respectively presented in the articles [30] and [10]. In [3], Beullens describes a new algorithm solving the trilinear form equivalence problem, harming the proposed parameters for ALTEQ. More recently, Narayanan, Qiao and Tang [20] presented an algorithm solving the same problem but also the matrix code equivalence problem in the case of $k$–dimensional spaces of $k\times k$ matrices. Their approach combines a collision list argument with a nice algebraic invariant and achieves a complexity in $\widetilde{\mathcal{O}}(q^{\frac{k}{2}})$. Finally, Ran and Samardjiska [24] designed an algorithm for the 3-tensor isomorphism problem which looks for triangles in tensor graphs. Such triangles exist in roughly $1/q$ of all instances of the problem. In these instances and for current parameters of MEDS and ALTEQ, their algorithm provides a speedup compared to all previous works.
@@ 1 The matrix code equivalence problem
Definition 1
Let $m,n,k$ be positive integers. An $m\times n$ matrix code of dimension $k$ is a $k$-dimensional $\mathbb{F}_{q}$-linear subspace of $\mathbb{F}_{q}^{m\times n}$.
@@
so any algorithm solving the case $m\leqslant n$ can also be used, after transposing the whole problem, to solve the case $n\leqslant m$. In the remainder of this article, we will always suppose that $m\leqslant n$. Moreover, we may switch $\mathcal{C}$ and its dual $\mathcal{C}^{\perp}$ in order to have $\dim(\mathcal{C}^{\perp})\leqslant\dim(\mathcal{C})$. Indeed, if $\mathcal{D}=P\mathcal{C}Q^{-1}$ then
$\mathcal{D}^{\perp}=(P^{-1})^{\top}\mathcal{C}^{\perp}Q^{\top}.$
Hence, the case where $mn-k=m=n$ may be reduced to an instance of $\mathsf{CMCE}$.
@@ 1.1.1 The trilinear forms equivalence problem
$f(Px,Qy,Rz)=g(x,y,z).$
The following well-known result shows the the matrix code equivalence problem reduces to the trilinear forms equivalence problem with the same parameters.
@@
$f$ $\colon(x,y,z)\mapsto\sum_{i,j,\ell}c_{ijk}x_{i}y_{j}z_{\ell}$
$g$ $\colon(x,y,z)\mapsto\sum_{i,j,\ell}d_{ijk}x_{i}y_{j}z_{\ell}.$
@@ 1.1.2 Alternate trilinear form equivalence problem
The alternate trilinear forms equivalence problem $\mathsf{ATFE}_{m,n,k}$ is the following. Given two alternate trilinear forms $f,g\colon\mathbb{F}_{q}^{m}\times\mathbb{F}_{q}^{m}\times\mathbb{F}_{q}^{k}\to\mathbb{F}_{q}$, find a matrix $P\in\operatorname{GL}_{n}(\mathbb{F}_{q})$ such that for any $x,y,z\in\mathbb{F}_{q}^{m}\times\mathbb{F}_{q}^{n}\times\mathbb{F}_{q}^{k}$,
@@
Rewriting an instance of $\mathsf{MCE}$ as an instance of $\mathsf{TFE}$ or of $3$–tensor isomorphism shows that the problem is symmetric in the three parameters $m,n,k$. In particular, we may choose to permute $m,n,k$ as we like in our algorithm in order to minimize its complexity. Moreover, as we will explain in Lemma 2, we may also switch $k$ for $mn-k$.
@@ 1.2 Related works on attacks on $\mathsf{MCE}$ and $\mathsf{ATFE}$
For the design of MEDS [10], the authors consider a graph-search based approach inspired from the works of Bouillaguet, Fouque and Véber [6] on the polynomial isomorphism problem for spaces of quadratic forms. This approach leads to an attack of complexity $\widetilde{\mathcal{O}}(q^{\frac{2}{3}(m+n)})$. Also, they consider the possibility of algebraic modelling which turns out to be harder than for $\mathsf{ATFE}$ since the unknown correspond to a pair of matrices $(P,Q)$ instead of a single one. They also study a “Leon–like” approach, a reference to Leon’s algorithm [17] for determining code equivalence that consists in harvesting minimum weight codewords to determine the code equivalence. When transposed to the matrix code setting, the Hamming weight is replaced by the rank and such an approach is nothing but the aforementioned MinRank based technique, which, following a recent result from Beullens [2] on Hamming metric code equivalence, can be combined with a collision search technique. MEDS’ parameter selection rests on the complexity of both algebraic attacks and Leon–like ones.
@@
First, Beullens proposed a graph–search–based technique to solve $\mathsf{ATFE}$. His attack turns out to be particularly efficient for small values of $n$. For instance, for $n$ odd, he could achieve a complexity in $\mathcal{O}(q^{{(n-5)}/{2}}n^{11}+q^{n-7}n^{6})$. This permitted to identify weak keys in [30].
@@
In [20], the authors introduced a new algorithm solving both $\mathsf{ATFE}$ and $\mathsf{MCE}$ in the cubic case $k=m=n$. We conclude this section by sketching the principle of this algorithm in order to point out the need for being in the cubic case.
@@
The idea of the algorithm consists first in guessing a pair $(x_{1},x_{1}^{\prime})\in{(\mathbb{F}_{q}^{n})}^{2}$ such that $x_{1}^{\prime}=Px_{1}$ and $f(x_{1},\cdot,\cdot)$ is a bilinear form of rank $n-1$. Next, due to the rank constraint, from $x_{1}$ can be deduced a unique $y_{1}$ (up to scalar multiplication) such that $f(x_{1},y_{1},\cdot)\equiv 0$. Similarly, they deduce a $z_{1}$ such that $f(\cdot,y_{1},z_{1})\equiv 0$, and an $x_{2}$ such that $f(x_{2},\cdot,z_{1})\equiv 0$ and so on. By this manner, they construct 3 sequences $x_{1},\dots,x_{n}$, $y_{1},\dots,y_{n}$ and $z_{1},\dots,z_{n}$ for $f$ and similarly construct $x_{1}^{\prime},\cdots,x_{n}^{\prime}$, $y_{1}^{\prime},\dots,y_{n}^{\prime}$ and $z_{1}^{\prime},\dots,z_{n}^{\prime}$ for $g$. Stacking these vectors as columns of $n\times n$ matrices, we get 6 matrices $X,Y,Z,X^{\prime},Y^{\prime}$ and $Z^{\prime}$ that will be invertible with a high probability.
@@
In view of this observation, the algorithm solving $\mathsf{MCE}$ consists in a collision search between two dictionaries. The first one collects pairs $(f_{x_{1}},x_{1})$, the left–hand term being the search key and the right–hand one being the corresponding value, and the second one collects pairs $(g_{x_{1}^{\prime}},x_{1}^{\prime})$. Once such a collision is found, determining the equivalence becomes easy (see [20] for further details). The space complexity of their algorithm is essentially the expected size of the computed dictionaries. Each dictionary has length $q^{(n-2)/2}$, and its entries are pairs consisting of a point of $\mathbb{P}^{n}(\mathbb{F}_{q})$ and a cubic 3-tensor of dimension $n\times n\times n$. Hence the total space complexity is $\mathcal{O}(n^{3}q^{(n-2)/2})$.
@@
It should be emphasized that the crux of their algorithm rests on the unique possibility (up to scalar multiplication) of passing from $x_{i}$ to $y_{i}$, from $y_{i}$ to $z_{i}$ and from $z_{i}$ to $x_{i+1}$. Such a technique is possible only because at each step, the corresponding bilinear form is represented by a rank $n-1$ matrix of size $n\times n$. Hence, their approach strongly rests on the fact that they lie in the cubic case $k=m=n$.
Ran and Samardjiska’s attack.
In [24], the authors describe a graph-based algorithm which solves the 3-tensor isomorphism problem in a specific case, namely when the graphs of the two isomorphic tensors contain cycles of length 3. This only happens in $\sim 1/q$ of all instances. However, contrary to [20], they are not limited to the cubic case. Their article contains a generic algorithm for the tensor isomorphism problem, and two versions adapted specifically to the $\mathsf{MCE}$ and $\mathsf{ATFE}$ problems. They all rely on the same modus operandi:
  • first, model by algebraic equations the existence of triangles in the graphs associated with the two tensors, and solve these using Gröbner basis techniques in order to construct lists of triangles found in each tensor’s graph;
  • then, for each pair of triangles in the two lists, try to construct the isometry by solving a system of linear and quadratic equations.
The authors provide timings showing that their attack is more efficient with real-world parameters than the previous ones, even going as far as breaking the designed Level I parameters of ALTEQ in under half an hour. However, this only applies to those $1/q$ of all instances in which the considered tensors do have triangles in their associated graphs.
Comparison with the present work.
We do not expect our work to beat the time complexity of [20] or [24] in the cases where they apply. However, our algorithm solves the $\mathsf{MCE}$ problem in a very broad range of parameters, and is not restricted to the cubic case like [20] or to those (very rare) tensors whose graphs contain triangles like [24]. Moreover, it improves upon the space complexity of [20] by a factor $\Theta(n^{2})$.